THE STRUCTURE, THE INTERNAL ROTATION OF N_{2} AND THE ROTATION OF THE ELECTRONIC TRANSITION MOMENT IN THE PARA-DIFLUOROBENZENE-N_{2} VAN DER WAALS COMPLEX

\underline{QUAN JU}, CHARLES S. PARMENTER, Department of Chemistry, Indiana University, Bloomington IN 47405.

Cold jet S_{1} ← S_{0} rotational band contours have been obtained at 0.3 cm^{−1} resolution for the p-difluorobenzene-N_{2} (pDFB-N_{2}) van der Waals complex and analyzed with an asymmetric top simulation program. Satisfactory simulation is possible only if it is assumed that the in-plane electronic transition moment of pDFB-N_{2} rotates about 37^{o} towards the F-F axis from its position normal to that axis in free pDFB. The vdW bond distances from the ring plane to the N_{2} bond parallel to that plane are 3.47Åand 3.42Åin the S_{1} and S_{0} states, respectively. Several transitions involving the internal rotation of N_{2} about an axis perpendicular to the ring plane have been assigned. The band positions are consistent with two-fold rotational barriers of 14.0 cm^{−1} and 19.0 cm^{−1} for the S_{0} and S_{1} states, respectively. These barriers are in agreement with those recently ^{a} determined from the assignments of different internal rotor transitions seen in the R2PI spectra.

aY. Hu, W. Lu and S. Yang, J. Photoch. Photobio. A, 106, 91(1997).